Find for each trinomial the value of k which makes the trinomial a perfect square.
y2 + 3y + k

Answer by fractalier(6550)

To find the k that makes this trinomial a perfect square, we take half of the 3 and then square it...so we get
k = [(1/2)(3)]^2
k = (3/2)^2
k = 9/4
thus we have
y^2 + 3y + 9/4 = (y + 3/2)^2